OpenStudy (help108):

MEDAL!! AND FAN What is the value of y in the equation 2(3y + 6 + 3) = 196 − 16?

alones (alones):

Intresitng problem 2(3y+6+3)=196−16 so simplify both sides of the equation. 2(3y+6+3)=196−16 6y+18=180 then Subtract 18 from both sides. 6y+18−18=180−18 6y=162 divide both sides by 6. 6y/6=162/6 y=27

rebeccaxhawaii (rebeccaxhawaii):

yes first distribute